using a 365-day calendar year, what is the interest due at closing for $54,750 mortgage loan at 8% interest for 23 days?
1
Expert's answer
2015-05-19T09:55:08-0400
If we use a 365-day calendar year, the interest due at closing for $54,750 mortgage loan at 8% interest for 23 days will be:
Interest payment IP = PV*i*t/365, where PV - present value of a loan, i - interest rate, t - number of days.
IP = 54,750*0.08*24/365 = $288